• Prove that (ab)n=anbn(ab)n=anbn is true for every natural number n
1
Expert's answer
2021-06-16T17:27:28-0400

Prove "(ab)^n=a^nb^n" is true for every natural number "n" by induction.

Let "P(n)" be the proposition that


"(ab)^n=a^nb^n"

for the natural number "n."

BASIS STEP:

"P(1)" ​is true because "(ab)^1=ab=a^1b^1."

This completes the basis step.


INDUCTIVE STEP: For the inductive hypothesis, we assume that "P(k)" is true for an arbitrary natural number "k." That is, we assume that


"(ab)^k=a^kb^k"

To carry out the inductive step using this assumption, we must show that when we assume that "P(k)" is true, then "P(k+1)" is also true. That is, we must show that


"(ab)^{k+1}=a^{k+1}b^{k+1}"

assuming the inductive hypothesis "P(k)." Under the assumption of "P(k)," we see that


"(ab)^k=a^kb^k"

"(ab)(ab)^k=(ab)a^kb^k"

"(ab)^{k+1}=aba^{k}b^{k}"

"(ab)^{k+1}=aa^{k}bb^{k}"

"(ab)^{k+1}=(aa^{k})(bb^{k})"




"(ab)^{k+1}=a^{k+1}b^{k+1}"

We have completed the inductive step.

Because we have completed the basis step and the inductive step, by mathematical induction we know that "P(n)" is true for all natural numbers "n."

That is, by principle of mathematical induction

"(ab)^n=a^nb^n"

for all natural numbers "n"